# A Differentiable Sparse Vector Quantization (SVQ) for Spatio-Temporal Forecasting

### v3

@@ 5.1 Benchmarks on various tasks
| PhyDNet(Guen & Thome, 2020) | 1614.7 | 0.9804 | 39.84 | 0.03709 | 2754.8 | 0.8615 | 23.26 | 0.32194 | 765.6 | 0.8322 | 23.41 | 0.50155 |
@@
| SimVP+SVQ (Frozen) | 1264.9 | 0.9851 | 34.07 | 0.02380 | 1408.6 | 0.9469 | 28.10 | 0.05535 | 364.6 | 0.9109 | 27.28 | 0.20988 |
| SimVP+SVQ (Learnable) | 1265.1 | 0.9851 | 34.06 | 0.02367 | 1414.9 | 0.9458 | 28.10 | 0.05776 | 360.2 | 0.9116 | 27.37 | 0.20658 |
@@
We explore both fixed (frozen) and learnable versions of SVQ on various forecasting tasks. Interestingly, we find that when codebook size is large, the performance of frozen, randomly-initialized codebook is on par with carefully learned codebook. Our comparison baselines consist of two categories: 1) Non-recurrent models including SimVP (Tan et al., 2022) and TAU (Tan et al., 2023a); 2) Recurrent-based models including ConvLSTM (SHI et al., 2015), PredNet (Lotter et al., 2017), PredRNN (Wang et al., 2017), PredRNN++ (Wang et al., 2018), MIM (Wang et al., 2019b), E3D-LSTM (Wang et al., 2019a), PhyDNet (Guen & Thome, 2020), MAU (Chang et al., 2021), PredRNNv2 (Wang et al., 2022b), and DMVFN (Hu et al., 2023). Baseline results are copied from the original OpenSTL paper (Tan et al., 2023b). To preclude ambiguity, we select the best MetaFormer of SimVP for each dataset, detailed in Appendix A.2.
@@
We present benchmark results of WeatherBench and three video prediction datasets (Human3.6M, KTH, and KittiCaltech) in Tables 1 and 2, respectively. Due to page limit, the benchmark results of TaxiBJ dataset are provided in Appendix B.4. These datasets have different characteristics. WeatherBench and TaxiBJ are macro forecasting tasks with low-frequency collection (30min or 1-6h). Human3.6M features subtle differences between frames as low-frequency signals. KittiCaltech poses challenges due to its rapidly changing backgrounds and limited training data. The KTH dataset, on the other hand, tests the capacity for long-horizon forecasting, requiring the prediction of 20 future frames based on 10 observed frames.
@@
However, despite the distinct characteristics among datasets, a common thread is the need for improved noise reduction coupled with enhanced representational capabilities, which can universally benefit their respective forecasting tasks. Notably, the SimVP+SVQ model achieves either the best or comparable performance across all datasets. For instance, on the WeatherBench-S temperature dataset, SVQ significantly improves the best baseline by 7.9% (1.105 → 1.018). On these three popular video prediction tasks, SVQ not only delivers a reduction in forecasting errors (average 9.4% decrease in MAE), but also significantly improves subjective image quality (average 17.3% decrease in LPIPS). Additional visualizations of forecasting samples can be found in Appendix D.
5.2 Boosting performance as a versatile plug-in
@@
In this section, SVQ serves as a versatile plug-in module applicable to various MetaFormers (Yu et al., 2022a). The adopted MetaFormers can be categorized into three types. 1) CNN-based: SimVPv1(IncepU) (Gao et al., 2022), SimVPv2(gSTA) (Tan et al., 2022), ConvMixer (Trockman & Kolter, 2023), ConvNeXt (Liu et al., 2022a), HorNet (Rao et al., 2022), and MogaNet (Li et al., 2022). 2) Transformer-based: ViT (Dosovitskiy et al., 2021), Swin Transformer (Liu et al., 2021), Uniformer (Li et al., 2023), Poolformer (Yu et al., 2022b), and VAN (Guo et al., 2023). 3) MLP-based: MLPMixer (Tolstikhin et al., 2021). We conduct experiments on WeatherBench-S temperature dataset because it is lightweight and fast for training.
@@
| ConvMixer(Trockman & Kolter, 2023) | 1.267 | 1.257 | 0.7073 | 0.6780 |
| ConvNeXt(Liu et al., 2022a) | 1.277 | 1.159 | 0.7220 | 0.6568 |
@@
5.3 Delicate balance between detail preservation and noise reduction
Investigating vector quantization’s (VQ) role in spatio-temporal forecasting, we evaluated several cutting-edge VQ methods akin to the SVQ framework, implemented as plug-in modules coordinated with the primary model. Table 4 shows that SVQ markedly improves forecasting as a plug-in, whereas other VQ approaches generally lead to heightened prediction errors. Enhanced detail retention within the representational capacity is linked to lower forecasting errors. Classic VQ methods suffer from notable information losses, as evidenced by a higher MSE of 1.854. In contrast, both residual VQ and grouped residual VQ outperform traditional VQ with lower MSEs of 1.213 and 1.174, respectively, affirming their ability to preserve intricate details.
It is commonly understood that the size of the codebook in clustering-based VQ is critical: larger codebooks capture more details, whereas smaller ones enhance noise reduction. To explore this trade-off, we analyzed how the codebook size influences the prediction MSE in grouped residual VQ. As Figure 4 indicates, MSE initially decreases but increases with overly large codebooks, echoing findings from (Yu et al., 2023b) that an excessively large codebook may degrade image generation performance. This underscores the necessity for dataset-specific tuning in clustering-based VQ approaches. In contrast, SVQ naturally balances detail preservation and noise reduction using sparse regression, negating the need for such fine-tuning.The codebook size in our approach exhibits a low-maintenance profile: performance remains solid across various codebook sizes, suggesting that utilizing a default large codebook can yield robust results without the need for extensive tuning. We are not suggesting that our SVQ outperforms other VQ methods in general image generation tasks, as this is beyond the scope of our current objective. Rather, we emphasize SVQ’s effectiveness as a noise reduction tool that directly enhances real-world spatio-temporal forecasting tasks, while the application to general image generation remains a topic for future exploration.
@@ 5.3 Delicate balance between detail preservation and noise reduction
| Residual VQ(Zeghidour et al., 2022) | 1.213 | 0.6910 |
| Grouped Residual VQ(Yang et al., 2023) | 1.174 | 0.6747 |
| Multi-headed VQ(Mama et al., 2021b) | 1.211 | 0.6994 |
| Residual VQ (Stochastic)(Lee et al., 2022) | 1.888 | 0.9237 |
| Residual Finite Scalar Quantization(Mama et al., 2021a) | 1.319 | 0.7505 |
@@
5.4 Robustness to artificial noise injection
To clearly demonstrate the noise reduction effect of SVQ, we conduct an experiment by adding artificial noise to the training data. Table 5 shows that the SVQ-equiped model experiences a much lower rise in MSE and MAE relative to the model without SVQ, regardless of the proportion of injected noise. For instance, when the proportion of injected noise is 10%, the MSE of the model without SVQ rises by 25.4%, whereas the model with SVQ shows a modest increase of just 3.2%. These results explicitly confirm that our method effectively mitigates noise by constraining latent patterns via quantization.
| Noise proportion | MSE | MAE | ||
| w/o SVQ | w SVQ | w/o SVQ | w SVQ | |
| =0 | 1.105 | 1.018 | 0.6567 | 0.6109 |
| =10% | 1.386(+25.4%) | 1.051(+3.2%) | 0.7702(+17.3%) | 0.6269(+2.6%) |
| =20% | 1.554(+40.7%) | 1.196(+17.5%) | 0.8282(+26.1%) | 0.6710(+9.8%) |
| =30% | 1.750(+58.4%) | 1.255(+23.2%) | 0.8821(+34.3%) | 0.6953(+13.8%) |
| =40% | 2.081(+88.3%) | 1.568(+54.0%) | 1.0031(+52.7%) | 0.7973(+30.5%) |
| =50% | 2.646(+139.4%) | 1.529(+50.2%) | 1.1339(+72.7%) | 0.7881(+29.0%) |
5.5 Ablation study
@@
We conduct a series of ablation studies on WeatherBench-S temperature dataset to understand the contribution of important designs based on the default setting: SVQ with 10000 codebook size, learnable codebook, and MAE loss.
Codebook size and learnability. We compare the effect of codebook size when codebook is learnable or frozen (Table 6). The results indicate that that increasing codebook size gradually improves performance. When significantly increasing codebook size to a large value, such as 10000, the performance difference between frozen and learnable codebook narrows, evidenced by a mere 0.5% (1.023 → 1.018). We observe that enlarging the codebook ensures comprehensive latent space coverage through random codes, negating the need for meticulous learning and allowing random initializations to rival learned counterparts. Additionally, our tailored SVQ structure surpasses alternative designs, including single-layer, bucket-shaped, and post-ReLU variants, while maintaining a similar parameter count.
| Learnability | Projection dim | Codebook size | MSE | MAE |
| Frozen | 128 | 10 | 1.070 | 0.6227 |
| 128 | 1000 | 1.044 | 0.6198 | |
| 128 | 10000 | 1.023 | 0.6131 | |
| Learnable | 128 | 10 | 1.060 | 0.6194 |
| 128 | 1000 | 1.048 | 0.6182 | |
| 128 | 10000 | 1.018 | 0.6109 | |
| 1280(Bucket-shape) | 1280 | 1.035 | 0.6149 | |
| None(One-layer) | 10000 | 1.043 | 0.6144 | |
| 128(Post-ReLU) | 10000 | 1.032 | 0.6136 |
@@
Self-learned sparse regression structure. The original SimVP model adopts MSE as prediction loss. We individually replace it with MAE loss and add the SVQ module. The results show that the joint use of SVQ and MAE loss is crucial for significantly improving the model’s performance (Table 7). We suggest that the sparsity of the weight matrix impacts vector representation learning and use kurtosis to quantify this after normalizing . Figure 5 demonstrates that both a learnable codebook and MAE loss contribute to increased sparsity. Analyzing four codebook initialization methods, in both learnable and fixed settings (Table 8), we find that a learnable codebook promotes sparsity irrespective of the initialization, indicating that sparsity is a self-learned property that enhances intermediate representation learning.
@@
6 Conclusions
In this work, we present Sparse regression-based Vector Quantization (SVQ), a concise yet effective method for spatiotemporal forecasting enhancement. SVQ elegantly balances detail preservation and noise reduction using sparse regression, realized via a two-layer MLP and a codebook, offering computational efficiency and full differentiability. Tested across diverse benchmarks, from weather to traffic and video prediction, SVQ consistently outperforms existing models, setting new performance standards without complex priors. Its differentiability and seamless integration with baseline models highlight SVQ as a significant advancement for efficient and effective spatio-temporal forecasting.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Supplemental Materials
The supplementary material for our work A Differentiable Sparse Vector Quantization (SVQ) for Spatio-Temporal Forecasting is organized as follows: Section A provides implementation details of SimVP model and VQ methods. Section B present extended quantitative results. Section C examines the effect of SVQ on latent representation. Finally, Section D shows additional qualitative results of forecasting samples and errors.
@@ Appendix A Implementation details
A.1 Dataset details
The details of the benchmark datasets are summarized as follows: WeatherBench (Rasp et al., 2020) and TaxiBJ (Zhang et al., 2017) are two macro forecasting tasks. These two datasets are collected at low frequencies (30min or 1-6h) and exhibit stable states, which is especially challenging for the model to capture subtle changes. Human3.6M (Ionescu et al., 2014), KittiCaltech (Geiger et al., 2013; Dollár et al., 2009), and KTH Action (Schüldt et al., 2004) are three popular video prediction tasks. Due to the diversity and complexity of real-world videos, these datasets have totally different characteristics. A summary of dataset statistics is provided in Table 9.
@@ A.1 Dataset details
A.2 Architecture configuration of SimVP
Table 10 reports the architectures of SimVP on all datasets. We select the best MetaFormer to replace the translator module based on OpenSTL benchmarks111https://openstl.readthedocs.io/en/latest/model_zoos/video_benchmarks.html222https://openstl.readthedocs.io/en/latest/model_zoos/weather_benchmarks.html333https://openstl.readthedocs.io/en/latest/model_zoos/traffic_benchmarks.html. The parameters remain unchanged, following the original configurations. It should be noted that, due to the reproducibility issue of ConvNeXt on the TaxiBJ dataset, we have opted to utilize gSTA as our backbone model.

### v4

@@ 4.1. Benchmarks on various forecasting tasks
| PhyDNet(Guen and Thome, 2020) | 1614.7 | 0.9804 | 39.84 | 0.03709 | 2754.8 | 0.8615 | 23.26 | 0.32194 | 765.6 | 0.8322 | 23.41 | 0.50155 |
@@
| SimVP+SVQ (Frozen codebook) | 1264.9 | 0.9851 | 34.07 | 0.02380 | 1408.6 | 0.9469 | 28.10 | 0.05535 | 364.6 | 0.9109 | 27.28 | 0.20988 |
| SimVP+SVQ (Learnable codebook) | 1265.1 | 0.9851 | 34.06 | 0.02367 | 1414.9 | 0.9458 | 28.10 | 0.05776 | 360.2 | 0.9116 | 27.37 | 0.20658 |
@@
The benchmark results of WeatherBench, TaxiBJ and three video prediction datasets (Human3.6M, KTH, and KittiCaltech) are presented in Tables 1, 2 and 3, respectively. These datasets have different characteristics. WeatherBench and TaxiBJ are macro forecasting tasks with low-frequency collection (30min or 1-6h). Human3.6M features subtle, low-frequency frame differences. KittiCaltech is challenging due to rapidly changing backgrounds and limited training data. The KTH dataset tests long-horizon forecasting, requiring the prediction of 20 future frames from 10 observed frames.
@@
Despite the distinct characteristics of each dataset, a consistent requirement across tasks is the need for effective noise reduction and pattern reservation, which universally benefit forecasting tasks. Notably, the SimVP+SVQ model achieves either the best or comparable performance across all datasets. For instance, on the WeatherBench-S temperature dataset, SVQ improves the best baseline by 7.9% (1.105 → 1.018). On three popular video prediction tasks, SVQ not only reduces forecasting errors (average 9.4% decrease in MAE), but also significantly improves subjective image quality (average 17.3% decrease in LPIPS).
Besides, we explore both fixed (frozen) and learnable versions of the SVQ codebook on these forecasting tasks. Interestingly, our findings reveal that with a large codebook size, the performance of a frozen, randomly-initialized codebook is on par with that of a carefully learned codebook. This observation aligns with our intuition: when allowed to choose a very large number of representative vectors to form a codebook, due to high orthogonality, a random choice is often as good as the one that is carefully chosen. Similar phenomena have been studied in the context of the column subset selection problem in matrix theory (Drineas et al., 2008; Deshpande and Rademacher, 2010).
To further investigate the impact of SVQ on latent representations, we compared the feature maps before and after applying SVQ. Using the KittiCaltech dataset as an example, Figure 4 demonstrates that the difference between foreground and background becomes more obvious after applying SVQ. In particular, road conditions and the sky are more clearly separated. These results suggest that SVQ effectively extracts meaningful patterns and enhances the discriminative capacity of latent representations, which in turn improves downstream forecasting performance.
4.2. Boosting performance as a versatile plug-in
@@
In this section, SVQ serves as a versatile plug-in module applicable to various MetaFormers (Yu et al., 2022a). The adopted MetaFormers include three types. 1) CNN-based: SimVPv1(IncepU) (Gao et al., 2022), SimVPv2(gSTA) (Tan et al., 2022), ConvMixer (Trockman and Kolter, 2023), ConvNeXt (Liu et al., 2022), HorNet (Rao et al., 2022), and MogaNet (Li et al., 2022). 2) Transformer-based: ViT (Dosovitskiy et al., 2021), Swin Transformer (Liu et al., 2021), Uniformer (Li et al., 2023), Poolformer (Yu et al., 2022b), and VAN (Guo et al., 2023). 3) MLP-based: MLPMixer (Tolstikhin et al., 2021). We conduct experiments on WeatherBench-S temperature dataset because it is lightweight and fast for training.
@@
| ConvMixer(Trockman and Kolter, 2023) | 1.267 | 1.257 | 0.7073 | 0.6780 |
| ConvNeXt(Liu et al., 2022) | 1.277 | 1.159 | 0.7220 | 0.6568 |
@@
4.3. Delicate balance between detail preservation and noise reduction
@@
| Residual VQ (RVQ) (Zeghidour et al., 2022) | 1.213 | 0.6910 |
| Grouped Residual VQ (GRVQ) (Yang et al., 2023) | 1.174 | 0.6747 |
| Multi-headed VQ (MHVQ) (Mama et al., 2021a) | 1.211 | 0.6994 |
| Stochastic Residual VQ (SRVQ) (Lee et al., 2022) | 1.888 | 0.9237 |
| Residual Finite Scalar Quantization (RFSQ) (Mama et al., 2021b) | 1.319 | 0.7505 |
@@
To investigate the role of VQ in spatio-temporal forecasting, we evaluated several state-of-the-art VQ methods that are structurally comparable to SVQ, each implemented as a plug-in module integrated with the backbone forecasting model. Table 5 shows that SVQ significantly improves forecasting performance, whereas other VQ methods lead to increased prediction errors. Enhanced detail retention within the representational capacity is associated with lower forecasting errors. In particular, traditional VQ suffers from notable information losses, as evidenced by a higher MSE of 1.854. In contrast, residual VQ and grouped residual VQ yield lower MSEs of 1.213 and 1.174, respectively, affirming their ability to preserve fine-grained information through recursive quantization.
It is commonly understood that the codebook size in clustering-based VQ is critical: larger codebooks capture more details, whereas smaller ones enhance noise reduction. To explore this trade-off, we compared how the codebook size influences the prediction MSE in Grouped Residual VQ (GRVQ) and SVQ. As Figure 5 indicates, the MSE of GRVQ initially decreases but increases with overly large codebooks, echoing findings from (Yu et al., 2023b) that an excessively large codebook may degrade image generation performance. This underscores the necessity for dataset-specific tuning in clustering-based VQ approaches. In contrast, SVQ naturally achieves a balance between preserving detail and reducing noise through sparse regression, eliminating the need for extensive fine-tuning. The codebook size in SVQ exhibits a low-maintenance profile: using a default large codebook can produce robust results without extensive tuning. We are not suggesting that our SVQ outperforms other VQ methods in general image generation tasks, as it is beyond the scope of our current objective. Rather, we emphasize SVQ’s effectiveness as a noise reduction tool that directly enhances real-world spatio-temporal forecasting tasks, while the application to general image generation remains a topic for future exploration.
4.4. Ablation study
@@
We conduct a series of ablation studies on WeatherBench-S temperature dataset to understand the contribution of key design components based on the default setting: SVQ with a codebook size of 10,000, learnable codebook, and MAE loss.
@@
Self-learned sparse regression structure. The original SimVP model adopts MSE as prediction loss. We individually replace it with MAE loss and add the SVQ module. As shown in Table 6, the joint use of SVQ and MAE loss is crucial for significantly improving the model’s performance. We suggest that the sparsity of the weight matrix impacts vector representation learning and use kurtosis to quantify this after normalizing . Figure 6 demonstrates that both a learnable codebook and MAE loss contribute to increased sparsity. Analyzing four codebook initialization methods in both learnable and fixed settings (Table 7), we find that a learnable codebook promotes sparsity irrespective of the initialization, indicating that sparsity is a self-learned property that enhances intermediate representation learning.
Codebook size and learnability. Table 8 compares the effects of codebook size—both learnable and frozen. Results show that increasing codebook size consistently enhances performance. However, when the size reaches 10,000, the performance gap between frozen and learnable codebooks narrows to just 0.5% (1.023 → 1.018). A larger codebook provides comprehensive coverage of the latent space through random codes, minimizing the need for meticulous learning. Consequently, models with randomly initialized codebooks perform similarly to those with learned ones. Additionally, our optimized SVQ structure outperforms alternative designs, including single-layer, bucket-shaped, and post-ReLU variants, while keeping a similar parameter count.
@@
| Learnability | Projection dim | Codebook size | MSE | MAE |
| Frozen | 128 | 10 | 1.070 | 0.6227 |
| 128 | 1000 | 1.044 | 0.6198 | |
| 128 | 10000 | 1.023 | 0.6131 | |
| Learnable | 128 | 10 | 1.060 | 0.6194 |
| 128 | 1000 | 1.048 | 0.6182 | |
| 128 | 10000 | 1.018 | 0.6109 | |
| 1280 (Bucket-shape) | 1280 | 1.035 | 0.6149 | |
| None (One-layer) | 10000 | 1.043 | 0.6144 | |
| 128 (Post-ReLU) | 10000 | 1.032 | 0.6136 |
Frozen module. The SVQ module consists of a two-layer MLP and a large codebook. The MLP can be seen as a projection from the input vector to the regression weights. We have already examined the impact of freezing the codebook on forecasting performance. To further investigate the effect of freezing the MLP, we conducted an ablation study in this section. The results, presented in Table 9, show that freezing the codebook only has a slight impact on forecasting performance, while freezing the MLP significantly impairs the performance. This highlights the critical role of the MLP in sparse regression, as it must be learned from data to generate the weights needed to combine codes from the codebook.
| Metric | Frozen module | |||
| None (All learnable) | Codebook | MLP | Both | |
| MSE | 1.018 | 1.023 | 1.060 | 1.093 |
| MAE | 0.6109 | 0.6131 | 0.6194 | 0.6387 |
4.5. Train Stability Issue
Although it is feasible to place the quantization module either before or after the translator, we found that for post-translator placement, the traditional VQ method (van den Oord et al., 2017) suffers pronounced instability and codebook collapse issues, as shown in Figure 7. It is essential to highlight that the backbones without VQ maintain their MSE within the acceptable range of approximately 1 to 2 (refer to Table 4). Yet, integrating traditional VQ causes a substantial rise in MSE values, exceeding 10 for different backbones—a level considered excessively high. Since two designs only differ in the placement order of quantization module and translator module, we hypothesize that this instability is attributed to the non-differentiability of the straight-through estimator, which introduces errors into the gradient flow for preceding modules. In contrast, our SVQ module never encounters such issues and remains highly stable throughout training. To maintain the integrity of all VQ methods, we opt for the pre-translator design in our main experiments, wherein quantization is executed before the translator module.
5. Conclusions
In this paper, we present Differentiable Sparse Soft-Vector Quantization (SVQ), an innovative and efficient technique for improving spatio-temporal forecasting. Unlike current state-of-the-art vector quantization methods, which often degrade performance, our approach is the first to show a beneficial impact on spatio-temporal forecasting tasks. SVQ elegantly tackles the inaccuracies in the optimization problem arising from non-differentiability and the restricted representational capabilities associated with hard-VQ. Tested across diverse benchmarks, from weather to traffic and video prediction, SVQ consistently outperforms pure baseline methods, setting new performance standards without complex prior models. Its differentiability and seamless integration with baseline models highlight SVQ as a significant advancement for efficient and effective spatio-temporal forecasting.
@@ Appendix A Implementation details
A.1. Dataset details
@@
A.2. Architecture configuration of SimVP
Table 11 reports the architectures of SimVP on all datasets. We select the best MetaFormer variant to replace the translator module based on OpenSTL benchmarks111https://openstl.readthedocs.io/en/latest/model_zoos/video_benchmarks.html222https://openstl.readthedocs.io/en/latest/model_zoos/weather_benchmarks.html333https://openstl.readthedocs.io/en/latest/model_zoos/traffic_benchmarks.html. The parameters remain unchanged, following the original configurations. It is noteworthy that due to reproducibility issues of ConvNeXt on the TaxiBJ dataset, we have opted to utilize gSTA as our backbone model.
A.3. Parameters of compared VQ methods
Table 5 presents a comparison of SVQ with several well-known VQ methods, reproduced using source code from the GitHub repository444https://github.com/lucidrains/vector-quantize-pytorch/tree/master. The parameters were kept consistent with the recommended settings to ensure performance, as detailed in Table 12. It should be noted that we found that increasing the codebook size for traditional VQ methods, such as Residual VQ and Multi-headed VQ, led to a considerable increase in GPU memory usage and extended the training time to impractical levels. This issue is one of the reasons these methods recommend adopting a default codebook size of 1024. To ensure fairness, we conducted an extensive experiment for VQ methods using the same codebook size (1024) in Appendix B.2.
Appendix B Additional quantitative results
B.1. Computational cost
Table 13 presents the computational costs of SVQ module and forecasting models. It shows that recurrent-based models have significantly higher FLOPs requirements, while non-recurrent models are more efficient. The proposed SVQ module is not only effective but also lightweight: across all datasets, it introduces only a minor increase in both parameter count and FLOPs. Notably, the overall computational cost of SimVP+SVQ remains substantially lower than that of recurrent-based models.
More importantly, the observed additional computational cost introduced by SVQ is primarily determined by the codebook size and dimensionality, and remains independent of the choice of base model. Consequently, if SVQ were applied to more complex baselines (e.g., ConvLSTM or PredRNN), the relative increase in FLOPs would be negligible.
B.2. Comprehensive comparison of VQ methods using the same codebook size
We extend Table 5 by setting the codebook size to the same value (1024) for compared VQ methods. They are comprehensively evaluated from different aspects: downstream performance (prediction MSE), codebook usage (perplexity), and computational complexity (FLOPS, inference FPS, and training time per epoch). Unlike other VQ methods that rely on a single code, our SVQ generates multiple regression weights to merge several codes. To evaluate its perplexity, we first normalize the regression weights and then convert them into binary form using a threshold set at times the standard deviation, where serves as the threshold value. Two thresholds (2 and 3) are utilized to obtain reasonable perplexity.
The quantitative and convergence results are shown in Table 14 and Figure 8. SVQ quickly converges to the lowest prediction error and satisfactory utilization of the codebook. Residual VQ with stochastic sampling has the highest codebook usage. However, its prediction MSE is worse than residual VQ without stochastic sampling. This demonstrates that forcibly improving codebook usage does not guarantee better downstream performance. SVQ generally outperforms the other VQ methods in computational efficiency, due to the approximation described in Section 3.1.